OpenStudy (anonymous):

A ball is thrown from an initial height of 4 feet with an initial upward velocity of 35ft/s. The ball's height h (in feet) after t seconds is given by the following. h= 4+35t-16t^2 Find all values of t for which the ball's height is feet.

OpenStudy (anonymous):

Find all values of t for which the ball's height is ___ feet ???

OpenStudy (anonymous):

sorry!! 22 feet

OpenStudy (anonymous):

h=22 4+35t-16t^2=22 -18+35t-16t^2=0 16t^2-35t+18=0

OpenStudy (anonymous):

Now, factor it and solve for t

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

how would you factor that? i used the quadratic formula and it has a radical in it

OpenStudy (anonymous):